OpenStudy (anonymous):

Evaluate and write your answer in scientific notation:(3 x 10^2)(1 x 10^-8)

OpenStudy (anonymous):

dose any one know how to do this

OpenStudy (zehanz):

Remember that is really 3 x 10^2 x 1 x 10^-8, so all multiplications. You may do multiplication in any order, so why not: (3 x 1) x (10^2 x10^-8) Calculate 3 x 1 and 10^2 x10^-8 seperately and you're done.

OpenStudy (anonymous):

can u solve it for me

OpenStudy (zehanz):

I will help you: First 3 x 1 = ... Second: 10^2 x 10^-8 = ...

OpenStudy (anonymous):

3*1=3

OpenStudy (zehanz):

For the second: it is of the form: \[a^b \cdot a^c=a^{b+c}\] So use this rule...

OpenStudy (anonymous):

i got the answere thank you

OpenStudy (zehanz):

YW!